# Emergent chirality and superfluidity of parity-doubled baryons in neutron stars

### v1

@@ II.3 Lagrangisns with the emergent chiral symmetries
III Mean-field analyses
@@ III.1 Nambu-Gor’kov formalism
by applying the Fierz transformations for the Dirac and color indices as summarized in Appendix C. Here we have defined $\psi_{C}=C\gamma^{0}\psi^{\ast}$ with $C=i\gamma^{2}\gamma^{0}$ where $\gamma^{\mu}=(\gamma^{0},\boldsymbol{\gamma})$ is the Dirac matrix. In Eq. (13), there are only the vector and axialvector channels and no other channels such as scalar and pseudoscalar channels. As a result, we find the possible quantum numbers of the pairings: the emergent-chiral-symmetric (${\cal S}$) and vector ($J^{P}=1^{-}$) pairings stemming from the bilinear form $\bar{\psi}_{C}\boldsymbol{\gamma}\gamma_{5}\tau_{a}\psi$ ($a=0,1,3$) and the emergent-chiral-antisymmetric (${\cal A}$) and axialvector ($J^{P}=1^{+}$) pairings from $\bar{\psi}_{C}\boldsymbol{\gamma}\tau_{2}\psi$ [^5] The condensates for the time-components, $\bar{\psi}_{C}\gamma_{0}\gamma_{5}\tau_{a}\psi$ and $\bar{\psi}_{C}\gamma_{0}\tau_{2}\psi$, are not considered in the present study, as explained in the text. Here $J^{P}$ indicates the spin-parity as a usual notation. We call the former the ${\cal S}_{a}1^{-}$ condensate, and the latter the ${\cal A}1^{+}$ condensate, as summarized in Table 2. Notice that the ${\cal S}_{a}1^{-}$ condensates are dependent on the directions in the $\mathrm{SU}(2)$ space ($a=0,1,3$). We comment that the $1^{-}$ pairing is denoted by ${}^{3}P_{1}$ and the $1^{+}$ pairing by ${}^{3}S_{1}$-${}^{3}D_{1}$ (mixing of $S$ and $D$ waves) when the nonrelativistic notation ${}^{2S+1}L_{J}$ with spin $S$ and angular momentum $L$ is adopted (cf. Fig. 1).
@@
Table 2: The quantum numbers for the pairings in the parity-doubled baryons are summarized. The quantum numbers for each pairing are given by ${\cal I}={\cal S}_{a}$ ($a=0,1,3$) or ${\cal A}$ indicating the representation of the emergent-chiral-symmetric or -antisymmetric representation and $J^{P}=1^{-}$ or $1^{+}$ for the spin-parity.
We classify the possible patterns of the stable pairings according to the signs of the coupling constants in the Lagrangian (13): (i) $g_{\perp}>0$ and $g_{\parallel}>0$, (ii) $g_{\perp}>0$ and $g_{\parallel}<0$, (iii) $g_{\perp}<0$ and $g_{\parallel}>0$, and (iv) $g_{\perp}<0$ and $g_{\parallel}<0$. When either of $g_{\perp}$ or $g_{\parallel}$ is negative, i.e., in the cases (ii), (iii), and (iv), we find that the time components of the interaction terms in Eq. (13) shift the chemical potential, and they play no important role for the pairing formations. Therefore, in the followings, we concentrate on the case (i).
In the case (i), we may neglect the repulsive parts in the interaction terms in Eq. (13) and leave the attractive parts only. Then, we simplify the Lagrangian as
@@ III.2 Phase diagram
The pairing formation in the superfluids induce the nonzero condensate for each $\boldsymbol{\Delta}_{a}$ and $\boldsymbol{\delta}$ in Eq. (17). As a simple situation, we suppose that the condensates $\boldsymbol{\Delta}_{a}$ and $\boldsymbol{\delta}$ are uniformly directed into the $z$ direction in the real space as constant vectors: $\boldsymbol{\Delta}_{a}=(0,0,\Delta_{a})$ for the ${\cal S}_{a}1^{-}$ condensate ($a=0,1,3$) and $\boldsymbol{\delta}=(0,0,\delta)$ for the ${\cal A}1^{+}$ condensate. Here $\Delta_{a}$ and $\delta$ are assumed to be constant values due to the uniformity. With this assumption, we obtain the single-particle energy for the (gapped) quasiparticles, stemming from the energy eigenvalues in the dressed propagator, $G(\{\boldsymbol{\Delta}_{a}\},\boldsymbol{\delta})$, in Eq. (20).
@@
Figure 4: The single-particle energies, ${\cal E}_{\boldsymbol{p}}^{+}$, ${\cal E}_{\boldsymbol{p}}^{-}$, ¯$\bar{\cal E}_{\boldsymbol{p}}^{+}$ and ¯$\bar{\cal E}_{\boldsymbol{p}}^{-}$, for the ${\cal A}1^{+}$ condensate with $\boldsymbol{\delta}=(0,0,\delta)$ are plotted. The values of the parameters are $\mu=1$ and $|\delta|=0.2$. The red surfaces (lines) at positive energy indicate ${\cal E}_{\boldsymbol{p}}^{+}$ and ¯$\bar{\cal E}_{\boldsymbol{p}}^{+}$, and the blue surfaces (lines) at negative energy indicate ${\cal E}_{\boldsymbol{p}}^{-}$ and ¯$\bar{\cal E}_{\boldsymbol{p}}^{-}$. (a) The energies are plotted on the $p_{x}$-$p_{z}$ plane with $p_{y}=0$. (b) The energies are plotted on the $p_{x}$ axis with $p_{z}=0$. (c) The energies for $p_{z}=\sqrt{\mu^{2}+|\delta|^{2}}$ are plotted on the $p_{x}$ axis with $p_{z}=0$. The dashed lines represent the approximate linear solutions in Eq. (55). (d) The energies are plot on the $p_{z}$ axis. Notice that, in (a), the cusps between the red and blue surfaces indicate the Dirac points at $\boldsymbol{p}_{\pm}^{\ast}$ given in Eq. (41). Notice also that, in (c), the red and blue dashed lines have the cusps (the discontinuity in the derivatives) at the Dirac points $p_{x}=0$ and $p_{z}=\sqrt{\mu^{2}+|\delta|^{2}}$, respectively, as indicated in Eq. (55).
@@
with $\xi=\Delta_{1}$ or $\delta$. The most stable states are obtained by minimizing each of $w_{\perp}(\Delta_{0})$, $w_{\perp}(\Delta_{3})$, $w_{\parallel}(\Delta_{1})$ and $w_{\parallel}(\delta)$ with respect to $\Delta_{0}$, $\Delta_{3}$, $\Delta_{1}$ and $\delta$, respectively. [^6] The possibility of the coexistence of different condensates, e.g., the coexistence of the ${\cal S}_{a}1^{-}$ and ${\cal A}1^{+}$ condensates, is not considered in the present study. Among them, the most stable states at zero temperature are determined by the relations among the coupling constants $g_{\perp}$ and $g_{\parallel}$:
@@
Figure 5: The schematic figure of the phase diagram is shown for the emergent chiral superfluids with the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}\times{\mathbb{Z}}_{2}$ emergent chiral symmetry [i.e., the $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}\times{\mathbb{Z}}_{2}$ full symmetry] and the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry [i.e., the $\mathrm{U}(2)_{\mathrm{L}}\times\mathrm{U}(2)_{\mathrm{R}}$ full symmetry]. Here e.c. sym. denotes the emergent chiral symmetry. In the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, the ${\cal S}_{0}1^{-}$ or ${\cal S}_{3}1^{-}$ condensate exists in the blue region ($g_{\perp}>g_{\parallel}$ and $g_{\perp}>0$), and the ${\cal S}_{1}1^{-}$ or ${\cal A}1^{+}$ condensate exists in the green region ($g_{\perp}<g_{\parallel}$ and $g_{\parallel}>0$). In the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry, the ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ condensates are degenerate on the red semiline ($g_{\perp}=g_{\parallel}>0$).
@@
Table 3: The condensates in the ground state are summarized for different emergent chiral symmetries and the coupling relations.
So far, we have considered the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}\times{\mathbb{Z}}_{2}$ emergent chiral symmetry by setting $g_{\perp}=g^{\prime}_{\perp}$. We may consider the original $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry by allowing $g_{\perp}$ and $g^{\prime}_{\perp}$ being different more generally. The condition $g_{\perp}\neq g^{\prime}_{\perp}$ would lead to resolving the degeneracy shown in the phase diagram in Fig. 5; either of the ${\cal S}_{0}1^{-}$ condensate and the ${\cal S}_{3}1^{-}$ condensate will be realized, and either of the ${\cal S}_{1}1^{-}$ condensate and the ${\cal A}1^{+}$ condensate will be realized. Similarly, on the semiline, only two degenerate condensates among the ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ condensates will be realized. In any case, because there exist emergent chiral superfluids in the ground state, our conclusion is not affected in essence.
@@ III.3 Dynamical symmetry breakings and their topology
for the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries, respectively. In all the cases, the baryon number symmetry $\mathrm{U}(1)_{\mathrm{L}+\mathrm{R}}$ is spontaneously broken, resulting in superfluidity. The chiral (axial) symmetry $\mathrm{U}(1)_{\mathrm{L}-\mathrm{R}}$ is unbroken. In contrast, the emergent chiral symmetries, $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$, are spontaneously broken to $\mathrm{U}(1)_{(1-2)(\mathrm{L}+\mathrm{R})}$ and $\mathrm{SU}(2)_{\mathrm{L}+\mathrm{R}}$, i.e., verctor-like emergent chiral symmetries interlocked by the left and right fermions, respectively. This implies the emergent chiral symmetry breaking. As for the spatial rotation, the $\mathrm{SO}(3)_{\rm S}$ rotational symmetry is broken to the $\mathrm{SO}(2)_{\rm S}$ symmetry, a rotation around the $z$-axis keeping the vector-type condensates such as $\boldsymbol{\Delta}_{a}=(0,0,\Delta_{a})$ for the ${\cal S}_{a}1^{-}$ condensate ($a=0,1,3$) and $\boldsymbol{\delta}=(0,0,\delta)$ for the ${\cal A}1^{+}$ condensate. The ${\mathbb{Z}}_{2({\rm L+R+S})}$ is defined as the action of a $\pi$ rotation of around the $x$ or $y$ axis represented by $\mathrm{diag}(1,-1,-1)$ or $\mathrm{diag}(-1,1,-1)\in\mathrm{SO}(3)_{\rm S}$ together with a $\pi$ phase rotation of the baryon symmetry $\boldsymbol{\Delta}_{a}\to e^{i\pi}\boldsymbol{\Delta}_{a},\boldsymbol{\delta}\to e^{i\pi}\boldsymbol{\delta}$ (or $\psi_{\mathrm{L}}\to e^{i\pi/2}\psi_{\mathrm{L}}$, $\psi_{\mathrm{R}}\to e^{i\pi/2}\psi_{\mathrm{R}}$ in terms of fermions degrees of freedom).777 This breaking pattern is the same with the polar phase of spin-1 spinor Bose-Einstein condensates [84]. Therefore, in such a way, the ${\cal S}_{a}1^{-}$ ($a=0,1,3$) or ${\cal A}1^{+}$ condensate induces a simultaneous dynamical breaking of the baryon number, emergent chiral and rotational symmetries.
@@
We leave some comments on the symmetry breakings in Eqs. (27) and (28). They indicate the new patterns of symmetry breaking at finite density. We may expect naively that the emergent chiral symmetry is usually restored at sufficiently high density. Indeed, the starting point of our discussion is based on the idea that the nucleon $N$ ($1/2^{+}$) and the excited nucleon $N^{\ast}(1535)$ ($1/2^{-}$) are degenerate at high density as the parity-doubled baryons. Contrary to this, the present study shows that the emergent chiral symmetry is dynamically broken by the pairings in the superfluids. It is also interesting that the vector-type condensate is realized in the superfluids. This is again different from the normal case that the scalar-type condensate is favored usually. Especially, the vector-type condensate cannot be seen in vacuum, because the realization of vector-type condensate is forbidden due to the Vafa-Witten theorem. [^8] The Vafa-Witten theorem indicates that that vector-like global symmetries cannot be spontaneously broken in the ground state. Thus, the emergent chiral superfluids exhibit unique features which are not seen in normal superfluids.
@@
for the $G_{\mathrm{L}}\times G_{\mathrm{R}}=\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$, and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries, respectively. The NG modes corresponding to these breakings are all of type-I with linear dispersion relations [85, 86, 87]. They may be called “emergent pions”, and consist of six quarks, with the baryon number two. These massless bosons can obtain masses to become pseudo-NG bosons [88] (like ordinary pions) when we add terms explicitly breaking the emergent chiral symmetries to the original Lagrangian. The naive and mirror assignments considered in the literature are such cases.
@@
where $S^{2}_{\rm S}\simeq\mathrm{SO}(3)_{\rm S}/\mathrm{SO}(2)_{\rm S}$. This OPS, $M_{\rm baryon,space}$, is the same with that of the polar phase of the spin-1 BEC, see the footnote 7, As mentioned, the spontaneous breaking of the baryon number symmetry $\mathrm{U}(1)_{\mathrm{L}+\mathrm{R}}$ is responsible for superfluidity. NG modes corresponding to this OPS are: a superfluid phonon as a type-I NG mode with a linear dispersion relation for the $\mathrm{U}(1)_{\mathrm{L}+\mathrm{R}}$ breaking, and magnons as a type-II NG mode with a quadratic dispersion relation for the $\mathrm{SO}(3)_{\rm S}$ symmetry breaking [87]. Unlike the NG bosons (emergent pions) for the emergent chiral symmetry breakings, these modes are exactly massless since the baryon and rotational symmetries are exact.
@@
Lastly, let us discuss topology and topological solitons associated with the dynamical symmetry breakings. The lower-dimensional homotopy groups $\pi_{0}$ to $\pi_{3}$ of $M_{\rm emergent\;chiral}$ are summarized in Table 4. Here 00 indicates a trivial group admitting no topological solitons. The $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ emergent chiral symmetry breaking admits axial strings supported by $\pi_{1}$.999 In the presence of explicit breaking terms, this axial string is attached by domain walls [89, 90]. The $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry breaking admits Skyrmions supported by $\pi_{3}$.

### v2

@@ II.3 Lagrangians with the emergent chiral symmetries
III Mean field analyses
@@ III.1 Nambu-Gor’kov formalism
by applying the Fierz transformations for the Dirac and color indices as summarized in Appendix C. Here we have defined $\psi_{C}=C\gamma^{0}\psi^{\ast}$ with $C=i\gamma^{2}\gamma^{0}$ where $\gamma^{\mu}=(\gamma^{0},\boldsymbol{\gamma})$ is the Dirac matrix. In Eq. (13), there are only the vector and axialvector channels and no other channels such as scalar and pseudoscalar channels. As a result, we find the possible quantum numbers of the pairings: the emergent-chiral-symmetric (${\cal S}$) and vector ($J^{P}=1^{-}$) pairings stemming from the bilinear form $\bar{\psi}_{C}\boldsymbol{\gamma}\gamma_{5}\tau_{a}\psi$ ($a=0,1,3$) and the emergent-chiral-antisymmetric (${\cal A}$) and axialvector ($J^{P}=1^{+}$) pairings from $\bar{\psi}_{C}\boldsymbol{\gamma}\tau_{2}\psi$ [^7] The condensates for the time components, $\bar{\psi}_{C}\gamma_{0}\gamma_{5}\tau_{a}\psi$ and $\bar{\psi}_{C}\gamma_{0}\tau_{2}\psi$, are not considered in the present study, as explained in the text. Here $J^{P}$ indicates the spin-parity as a usual notation. We call the former the ${\cal S}_{a}1^{-}$ condensate and the latter the ${\cal A}1^{+}$ condensate, as summarized in Table 2. Notice that the ${\cal S}_{a}1^{-}$ condensates are dependent on the directions in the $\mathrm{SU}(2)$ space ($a=0,1,3$). We comment that the $1^{-}$ pairing is denoted by ${}^{3}P_{1}$ and the $1^{+}$ pairing by ${}^{3}S_{1}$-${}^{3}D_{1}$ (mixing of $S$ and $D$ waves) when the nonrelativistic notation ${}^{2S+1}L_{J}$ with spin $S$ and angular momentum $L$ is adopted (cf. Fig. 1).
@@
Table 2: Quantum numbers for the pairings in the parity-doubled baryons are summarized. The quantum numbers for each pairing are given by ${\cal I}={\cal S}_{a}$ ($a=0,1,3$) or ${\cal A}$ indicating the representation of the emergent-chiral-symmetric or -antisymmetric representation and $J^{P}=1^{-}$ or $1^{+}$ for the spin-parity.
We classify the possible patterns of the stable pairings according to the signs of the coupling constants in the Lagrangian (13): (i) $g_{\perp}>0$ and $g_{\parallel}>0$, (ii) $g_{\perp}>0$ and $g_{\parallel}<0$, (iii) $g_{\perp}<0$ and $g_{\parallel}>0$, and (iv) $g_{\perp}<0$ and $g_{\parallel}<0$. When either $g_{\perp}$ or $g_{\parallel}$ is negative, i.e., in the cases (ii), (iii), and (iv), we find that the time components of the interaction terms in Eq. (13) shift the chemical potential, and they play no important role for the pairing formations. Therefore, in the following, we concentrate on case (i).
In case (i), we may neglect the repulsive parts in the interaction terms in Eq. (13) and leave the attractive parts only. Then, we simplify the Lagrangian as
@@ III.2 Phase diagram
The pairing formation in the superfluids induce the nonzero condensate for each $\boldsymbol{\Delta}_{a}$ and $\boldsymbol{\delta}$ in Eq. (17). As a simple situation, we suppose that the condensates $\boldsymbol{\Delta}_{a}$ and $\boldsymbol{\delta}$ are uniformly directed into the $z$ direction in the real space as constant vectors: $\boldsymbol{\Delta}_{a}=(0,0,\Delta_{a})$ for the ${\cal S}_{a}1^{-}$ condensate ($a=0,1,3$) and $\boldsymbol{\delta}=(0,0,\delta)$ for the ${\cal A}1^{+}$ condensate. Here, $\Delta_{a}$ and $\delta$ are assumed to be constant values due to the uniformity. With this assumption, we obtain the single-particle energy for the (gapped) quasiparticles, stemming from the energy eigenvalues in the dressed propagator, $G(\{\boldsymbol{\Delta}_{a}\},\boldsymbol{\delta})$, in Eq. (20).
@@
Figure 4: Single-particle energies, ${\cal E}_{\boldsymbol{p}}^{+}$, ${\cal E}_{\boldsymbol{p}}^{-}$, ¯$\bar{\cal E}_{\boldsymbol{p}}^{+}$ and ¯$\bar{\cal E}_{\boldsymbol{p}}^{-}$, for the ${\cal A}1^{+}$ condensate with $\boldsymbol{\delta}=(0,0,\delta)$ are plotted. The values of the parameters are $\mu=1$ and $|\delta|=0.2$. The red surfaces (lines) at positive energy indicate ${\cal E}_{\boldsymbol{p}}^{+}$ and ${\cal E}_{\boldsymbol{p}}^{-}$, and the blue surfaces (lines) at negative energy indicate ¯$\bar{\cal E}_{\boldsymbol{p}}^{+}$ and ¯$\bar{\cal E}_{\boldsymbol{p}}^{-}$. (a) The energies are plotted on the $p_{x}$-$p_{z}$ plane with $p_{y}=0$. (b) The energies are plotted on the $p_{x}$ axis with $p_{z}=0$. (c) The energies for $p_{z}=\pm\sqrt{\mu^{2}+|\delta|^{2}}$ are plotted on the $p_{x}$ axis with $p_{y}=0$. The dashed lines represent the approximate linear solutions in Eq. (55). (d) The energies are plot on the $p_{z}$ axis. Notice that, in (a), the cusps between the red and blue surfaces indicate the Dirac points at $\boldsymbol{p}_{\pm}^{\ast}$ given in Eq. (41). Notice also that, in (c), the red and blue dashed lines have the cusps (the discontinuity in the derivatives) at the Dirac points $p_{x}=0$ and $p_{z}=\pm\sqrt{\mu^{2}+|\delta|^{2}}$, respectively, as indicated in Eq. (55).
@@
with $\xi=\Delta_{1}$ or $\delta$. The most stable states are obtained by minimizing each of $w_{\perp}(\Delta_{0})$, $w_{\perp}(\Delta_{3})$, $w_{\parallel}(\Delta_{1})$ and $w_{\parallel}(\delta)$ with respect to $\Delta_{0}$, $\Delta_{3}$, $\Delta_{1}$ and $\delta$, respectively.  [^8] The possibility of the coexistence of different condensates, e.g., the coexistence of the ${\cal S}_{a}1^{-}$ and ${\cal A}1^{+}$ condensates, is not considered in the present study. Among them, the most stable states at zero temperature are determined by the relations among the coupling constants $g_{\perp}$ and $g_{\parallel}$
@@
Figure 5: Schematic figure of the phase diagram is shown for the emergent chiral superfluids with the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}\times{\mathbb{Z}}_{2}$ emergent chiral symmetry [i.e., the $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}\times{\mathbb{Z}}_{2}$ full symmetry] and the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry [i.e., the $\mathrm{U}(2)_{\mathrm{L}}\times\mathrm{U}(2)_{\mathrm{R}}$ full symmetry]. Here e.c. sym. denotes the emergent chiral symmetry. In the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, the ${\cal S}_{0}1^{-}$ or ${\cal S}_{3}1^{-}$ condensate exists in the blue region ($g_{\perp}>g_{\parallel}$ and $g_{\perp}>0$), and the ${\cal S}_{1}1^{-}$ or ${\cal A}1^{+}$ condensate exists in the green region ($g_{\perp}<g_{\parallel}$ and $g_{\parallel}>0$). In the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry, the ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$, and ${\cal A}1^{+}$ condensates are degenerate on the red semiline ($g_{\perp}=g_{\parallel}>0$).
@@
Table 3: Condensates in the ground state are summarized for different emergent chiral symmetries and the coupling relations.
So far, we have considered the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}\times{\mathbb{Z}}_{2}$ emergent chiral symmetry by setting $g_{\perp}=g^{\prime}_{\perp}$. We may consider the original $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry by allowing $g_{\perp}$ and $g^{\prime}_{\perp}$ to be different more generally. The condition $g_{\perp}\neq g^{\prime}_{\perp}$ would lead to resolving the degeneracy shown in the phase diagram in Fig. 5; either the ${\cal S}_{0}1^{-}$ condensate or the ${\cal S}_{3}1^{-}$ condensate will be realized, and either the ${\cal S}_{1}1^{-}$ condensate or the ${\cal A}1^{+}$ condensate will be realized. Similarly, on the semiline, only two degenerate condensates among the ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ condensates will be realized. In any case, because there exist emergent chiral superfluids in the ground state, our conclusion is not affected in essence.
@@ III.3 Dynamical symmetry breakings and their topology
for the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries, respectively. In all the cases, the baryon number symmetry $\mathrm{U}(1)_{\mathrm{L}+\mathrm{R}}$ is spontaneously broken, resulting in superfluidity. The chiral (axial) symmetry $\mathrm{U}(1)_{\mathrm{L}-\mathrm{R}}$ is unbroken. In contrast, the emergent chiral symmetries, $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$, are spontaneously broken to $\mathrm{U}(1)_{(1-2)(\mathrm{L}+\mathrm{R})}$ and $\mathrm{SU}(2)_{\mathrm{L}+\mathrm{R}}$, i.e., vector-like emergent chiral symmetries interlocked by the left and right fermions, respectively. This implies the emergent chiral symmetry breaking. As for the spatial rotation, the $\mathrm{SO}(3)_{\rm S}$ rotational symmetry is broken to the $\mathrm{SO}(2)_{\rm S}$ symmetry, a rotation around the $z$-axis keeping the vector-type condensates such as $\boldsymbol{\Delta}_{a}=(0,0,\Delta_{a})$ for the ${\cal S}_{a}1^{-}$ condensate ($a=0,1,3$) and $\boldsymbol{\delta}=(0,0,\delta)$ for the ${\cal A}1^{+}$ condensate. The ${\mathbb{Z}}_{2({\rm L+R+S})}$ is defined as the action of a $\pi$ rotation around the $x$ or $y$ axis represented by $\mathrm{diag}(1,-1,-1)$ or $\mathrm{diag}(-1,1,-1)\in\mathrm{SO}(3)_{\rm S}$ together with a $\pi$ phase rotation of the baryon symmetry $\boldsymbol{\Delta}_{a}\to e^{i\pi}\boldsymbol{\Delta}_{a},\boldsymbol{\delta}\to e^{i\pi}\boldsymbol{\delta}$ (or $\psi_{\mathrm{L}}\to e^{i\pi/2}\psi_{\mathrm{L}}$, $\psi_{\mathrm{R}}\to e^{i\pi/2}\psi_{\mathrm{R}}$ in terms of fermions degrees of freedom). 999 This breaking pattern is the same with the polar phase of spin-1 spinor Bose-Einstein condensates [87]. Therefore, in such a way, the ${\cal S}_{a}1^{-}$ ($a=0,1,3$) or ${\cal A}1^{+}$ condensate induces a simultaneous dynamical breaking of the baryon number, emergent chiral and rotational symmetries.
@@
We leave some comments on the symmetry breakings in Eqs. (27) and (28). They indicate the new patterns of symmetry breaking at finite density. We may expect naively that the emergent chiral symmetry is usually restored at sufficiently high density. Indeed, the starting point of our discussion is based on the idea that the nucleon $N$ ($1/2^{+}$) and the excited nucleon $N^{\ast}(1535)$ ($1/2^{-}$) are degenerate at high density as the parity-doubled baryons. Contrary to this, the present study shows that the emergent chiral symmetry is dynamically broken by the pairings in the superfluids. It is also interesting that the vector-type condensate is realized in the superfluids. This is again different from the normal case that the scalar-type condensate is favored usually. Especially, the vector-type condensate cannot be seen in vacuum, because the realization of vector-type condensate is forbidden due to the Vafa-Witten theorem.  [^10] The Vafa-Witten theorem indicates that that vector-like global symmetries cannot be spontaneously broken in the ground state [88, 89]. Thus, the emergent chiral superfluids exhibit unique features which are not seen in normal superfluids.
@@
for the $G_{\mathrm{L}}\times G_{\mathrm{R}}=\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries, respectively. The NG modes corresponding to these breakings are all of type-I with linear dispersion relations [90, 91, 92]. They may be called “emergent pions” and consist of six quarks, with the baryon number two. These massless bosons can obtain masses to become pseudo-NG bosons [93] (like ordinary pions) when we add terms explicitly breaking the emergent chiral symmetries to the original Lagrangian. The naive and mirror assignments considered in the literature are such cases.
@@
where $S^{2}_{\rm S}\simeq\mathrm{SO}(3)_{\rm S}/\mathrm{SO}(2)_{\rm S}$. This OPS, $M_{\rm baryon,space}$, is the same with that of the polar phase of the spin-1 Bose-Einstein condensate (BEC); see footnote 9. As mentioned, the spontaneous breaking of the baryon number symmetry $\mathrm{U}(1)_{\mathrm{L}+\mathrm{R}}$ is responsible for superfluidity. NG modes corresponding to this OPS are a superfluid phonon as a type-I NG mode with a linear dispersion relation for the $\mathrm{U}(1)_{\mathrm{L}+\mathrm{R}}$ breaking and magnons as a type-II NG mode with a quadratic dispersion relation for the $\mathrm{SO}(3)_{\rm S}$ symmetry breaking [92]. Unlike the NG bosons (emergent pions) for the emergent chiral symmetry breakings, these modes are exactly massless since the baryon and rotational symmetries are exact.
@@
Lastly, let us discuss topology and topological solitons associated with the dynamical symmetry breakings. The lower-dimensional homotopy groups $\pi_{0}$ to $\pi_{3}$ of $M_{\rm emergent\;chiral}$ are summarized in Table 4. Here 00 indicates a trivial group admitting no topological solitons. The $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ emergent chiral symmetry breaking admits axial strings supported by $\pi_{1}$111111 In the presence of explicit breaking terms, this axial string is attached by domain walls [94, 95]. The $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry breaking admits Skyrmions supported by $\pi_{3}$.